OpenStudy (anonymous):

It is known that a bag of apples weighing 5.23 kg also weighs 11.53 pounds. From this information, derive an equation converting kg to pounds. Use it to calculate how much would a bag of oranges weighing 4.87 kg would weigh in pounds.

OpenStudy (id21):

5.23 kg=11.53 pounds Lets find how much pounds is 1 kg: \[\frac{5.23}{5.23}kg=\frac{11.53}{5.23}pounds\] so \[1kg \approx 2.205pounds\] Then 4.87 kg will be \[4.87\times 2.205\approx 10.738 pounds\]

OpenStudy (anonymous):

The required equation is y = 2.2x (x is kg, y is pounds). When x is 4.87, y is 2.2* 4.87 = 10.71 pounds approx.